monochromatic beam of light of wavelength 663nm is incident normal to the surface produces a pressure of 10^-6 N/m^2. the surface absorbs 75% of light and reflects the rest. find the number of photons per unit area per second:

Grade:12th pass

1 Answers

mahima pareek
5 years ago
The expression for pressure exerted when there is reflection and absorption is 
        P = I(1 + r) /c
Where P = pressure exerted 
             I = intensity of the incident light
             r = reflection coefficient 
            c = speed of light
Here r = 100 - 75/100 = 0.25
Now we know 
           I = nE/Atc 
Where.   n = no.  of photons incident 
               E = energy of each photon 
                A = area of the metal plate 
                t = time of incidence of photons 
Now we are given P and asked about n/At
So our formula for pressure becomes 
         P =( The thing which is asked) E(1 + 0.25)/c
Now n/At = Pc/E(1.25)
                  Answer is probably 9.84 * 10 to the power 20
 
PS i did not solve it as I don't have special characters in phone 
Approve if the asnwer is correct